How Many ounces of a 40% alchol solution must be mixed with 8 ounces of 45% alchol solution to make 44%
Answer by checkley77(12844)

.45*8+.40X=.44(8+X)
3.6+.40X=3.52+.44X
.40X-.44X=3.52-3.6
-.04X=-.08
X=-.08/-.04
X=2 OZ. OF 45% SOLUTION IS USED.
8-2=6 OZ. OF 40% SOLUTION IS USED.
PROOF:
.45*8+.40*2=.44(8+2)
3.6+.8=.44*10
4.4=4.4